# Assessing High-Risk Systems: An EU AI Act Verification Framework

### v1

@@ 4.1. R1: Human Agency and Oversight
Article 14 of the AI Act establishes the principle of human agency and oversight as a cornerstone of trustworthy AI. It requires that AI systems be designed and implemented so that they operate under meaningful human control and that ultimate responsibility remains with human operators. In practice, this means that individuals supervising the system must understand its capabilities and limitations, be able to intervene when necessary, and override its decisions to prevent or mitigate risks. The article also reinforces user autonomy by mandating transparency, informed consent, and the possibility to opt out of AI-driven functionalities. Together, these obligations ensure that human judgment and accountability remain central throughout the design, deployment, and use of AI systems.
Requirement Guidelines Methods Type Target
R1.1: Human in the loop AIA (Art. 14.4.d–e) M1.1: Manual override — Confirm that a human operator can take control of the AI system at any time during operation. C P, FP
AIA (Art. 14.4.a–c) M1.2: Supervisory control — Assess whether human operators are able to supervise the AI system and intervene when necessary. C P, FP
@@
R1.2: User Autonomy AIA (Art. 50.1) M1.4: Informed consent — Check that users are provided with clear information about AI usage and are able to give or withhold consent. C FP
ISACA manual (Sec. 3.2, p.137) M1.5: User preferences — Review whether the system allows adaptation of its behaviour based on user-set preferences. C FP
ISACA manual (Sec. 3.2, p.137) M1.6: Opt-in/out — Confirm that users can enable or disable AI functionalities according to their choices. C FP
R1.3: Oversight Mechanisms ISACA manual (Sec. 2.11.1) M1.7: Audit trails — Ensure that system actions and decisions are recorded in tamper-evident logs for traceability. C P
@@ 4.2. R2: Technical Robustness and Safety
Requirement Guidelines Methods Type Target
R2.1: Resilience & Reliability ISACA manual (Sec. 1.4.4) M2.1: Stress testing — Assess whether the system has been exposed to extreme loads or edge cases to identify potential failure points. T P
ISACA manual (Sec. 2.8.6) M2.2: Drift detection — Confirm that mechanisms are in place to monitor data distribution changes and evaluate behaviour under such shifts. T D, P
@@
AIA (Art. 15.4); ISACA manual (Sec. 2.9.4) M2.4: Fail-over architecture — Ensure redundant components exist to maintain functionality in case of failure. C P, FP
R2.2: Robustness to Attacks Brundage et al. (Brundage2020); ISACA manual (Sec. 2.13.3) M2.5: Red teaming — Examine whether adversarial testing is conducted to uncover vulnerabilities. T P, FP
R2.3: Fallback & Fail-safe HLEG Guidelines; AIA (Art. 14.4.e) M2.6: Human fallback — Check that the system includes a mechanism to transfer control to a human in the event of failure. C P, FP
R2.4: Accuracy & Uncertainty ISACA manual (Sec. 1.4.3) M2.7: Model calibration — Review whether confidence scores are appropriately aligned with model performance. T M
@@ 4.3. R3: Privacy and Data Governance
Article 10 of the AI Act establishes specific obligations concerning data governance and data quality for high-risk AI systems, requiring that datasets be relevant, representative, free of errors, and complete to the extent possible. These provisions complement the data protection principles set out in the GDPR, particularly Articles 5, 25, and 32, which enshrine purpose limitation, privacy by design, and security of processing. Together, they ensure that data is collected, processed, and stored in a lawful, fair, and transparent manner. The combined framework mandates both organisational and technical measures—such as encryption, pseudonymisation, and role-based access control—to prevent unauthorised use and ensure accountability for data-related decisions. Effective data governance thus underpins not only compliance with legal obligations but also the technical robustness and trustworthiness of AI systems across their lifecycle.
Requirement Guidelines Methods Type Target
@@
GDPR (Art. 32.1.a); ISACA manual (Sec. 2.7.3) M3.5: Encryption — Check that data is encrypted in storage and transmission. T D, P
R3.4: Data Provenance ISACA manual (Sec. 2.7.5) M3.6: Lineage tracking — Confirm that data transformations are logged to ensure traceability. C D, P
ISACA manual (Sec. 3.1.2) M3.7: Metadata versioning — Verify dataset versions and schema changes are tracked. C D, P
@@ 4.4. R4: Transparency
Transparency is one of the most fundamental requirements for trustworthy and accountable AI. Under the AI Act, AI systems must be designed and documented in a way that allows users, regulators, and auditors to understand their capabilities, limitations, and decision-making logic. Transparency encompasses both technical and communicative aspects: it includes model explainability, data and process traceability, disclosure of AI usage to end users, and interpretability of outputs in accessible language. Together, these mechanisms ensure that AI operations remain observable and interpretable across all stages of their lifecycle, thereby fostering accountability and enabling meaningful human oversight.
Requirement Guidelines Methods Type Target
R4.1: Explainability ISACA manual (Sec. 3.2, p.139) M4.1: Local interpretability — Verify that the system provides human-understandable explanations of model outputs using recognised local explanation techniques (e.g., feature attribution or surrogate models). T D, M, FP
Mundhenk et al. (mundhenk2020) M4.2: Saliency maps — Check if visualization tools highlight relevant input regions. T D, M, FP
R4.2: Model & Data Traceability GPAI CoP (Sec. Transparency) M4.3: Model cards — Ensure that standardized documentation is available and maintained. C M, P
@@
R4.3: Disclosure of AI Use AIA (Art. 50.1) M4.5: AI identity notices — Verify that users are clearly informed when interacting with an AI system. C FP
GPAI CoP; ISACA manual (Sec. 3.2, p.139) M4.6: Disclaimers — Check whether limitations and assumptions are transparently disclosed. C FP
AIA (Art. 13.3.b.III) M4.7: Risk communication — Assess whether the system communicates known risks and uncertainties. C FP
R4.4: Interpretability for End Users AIA (Art. 13.2) M4.8: Understandable language — Confirm that technical terms are explained using plain language. C FP
@@ 4.5. R5: Diversity, Non-discrimination and Fairness
Diversity, non-discrimination, and fairness are central to ensuring that AI systems respect fundamental rights and avoid reinforcing structural biases. Articles 9, 10, and 27 of the AI Act require that systems be designed, trained, and deployed in a way that prevents discriminatory outcomes and ensures inclusive access. This includes assessing data representativeness, applying fairness metrics across protected groups, and incorporating participatory design practices that reflect societal diversity. Fairness thus spans both technical and organisational domains, linking data quality, model evaluation, and stakeholder engagement to the ethical and legal accountability of AI systems.
Requirement Guidelines Methods Type Target
R5.1: Fairness in Data ISACA manual (Sec. 3.2, p.138) M5.1: Bias detection — Evaluate whether datasets are examined for representation gaps. T D, P
ISACA manual (Sec. 3.2, p.138) M5.2: Sampling strategies — Confirm that sampling ensures balanced representation. C D, P
R5.2: Fairness in Algorithms ISACA manual (Sec. 3.2, p.138) M5.3: Protected-class fairness metrics — Verify that fairness metrics are applied across protected groups. T D, M, P
R5.3: Inclusive Design AIA (Rec. 132) M5.4: Accessibility — Check that the system supports assistive technologies. C FP
R5.4: Stakeholder Engagement AIA (Rec. 165) M5.5: Participatory design — Assess whether affected users and communities are involved in design and testing. C FP, P
@@ 4.6. R6: Societal and Environmental Well-being
Requirement Guidelines Methods Type Target
R6.1: Environmental Impact UNESCO (Rec. 84); ISACA manual (Sec. 1.17.7) M6.1: Energy-use tracking — Verify that the system monitors computational energy consumption and carbon footprint. T P, FP
R6.2: Ethical Alignment AIA (Art. 27.1) M6.2: Rights-based system — Assess whether system design aligns with fundamental rights. C P
@@ 4.7. R7: Accountability
Requirement Guidelines Methods Type Target
R7.1: Responsibility Assignment ISACA manual (Sec. 3.2, p.129) M7.1: RASCI charts — Confirm that responsibility matrices define roles for each lifecycle activity. C P
R7.2: Auditability ISO/IEC 42001:2023 (Ann.B.6.2.8); ISACA manual (Sec. 2.11.1) M7.2: Logging — Verify that detailed, tamper-evident logs of system operations are maintained and reviewed. C P
@@
R7.4: Documentation Integrity ISO/IEC 42001:2023 (Ann.B.6.2.5); ISACA manual (Sec. 2.11.1) M7.5: Version-controlled records — Assess whether documentation is versioned and traceable over time. C P
@@ 4.8. R8: Quality Management
Article 17 of the AI Act requires providers of high-risk AI systems to establish, implement, document, and maintain a quality management system (QMS). The QMS ensures that the organisation applies consistent procedures for design, testing, data management, and post-market monitoring. It provides a structured approach for achieving and maintaining conformity with the requirements of the AI Act, aligning with recognised standards such as ISO 9001 and ISO/IEC 42001. Quality management thereby supports traceability, continuous improvement, and long-term compliance across the AI system lifecycle.
Requirement Guidelines Methods Type Target
@@ 4.9. R9: Risk Management
Article 9 of the AI Act establishes the obligation to implement a continuous and systematic risk management process throughout the lifecycle of high-risk AI systems. This process covers the identification, analysis, evaluation, and control of risks, as well as post-market monitoring and review. It must ensure that residual risks are acceptable relative to the intended use and that appropriate mitigation measures are applied. The methodology used here aligns with a set of horizontal, cross-sectoral standards, including ISO/IEC 23894 (AI risk management), ISO/IEC 42001 (AI management system), ISO 31000 and ISO 31010 (General risk management principles and techniques), as well as data- and robustness-related standards such as ISO/IEC 5259, ISO 8000, ISO/IEC TR 24027, ISO/IEC 24029, and ISO/IEC 24970. In coming days, when harmonized standards under the AI Act, particularly prEN 18228 (AI Risk Management) and prEN 18286 (AI Quality Management System) will be finalized, they will provide the solid structure for conformity assessment and presumption of conformity once a system is categorized as high-risk.
@@
However, it is essential to underline that none of these standards, both international ISO/IEC standards and future harmonized European standards, determine whether an AI system is classified as high-risk under Article 6. Classification is governed exclusively by the legal criteria of the AI Act, particularly Annex I and Annex III. Standards are used in this report solely to support structured analysis, documentation and evidence generation, and they become fully applicable only after an AI system has been classified as high-risk.
In addition, the methodology deliberately excludes sector-specific or application-specific standards, such as functional-safety standards (e.g., IEC 61508 or ISO 13849), medical-device standards (e.g., ISO 14971) or automotive safety standards (e.g., ISO 26262). These frameworks are highly relevant within their respective regulatory regimes but are not part of the horizontal, cross-sector risk-classification process defined by the AI Act and therefore are not included in the present analysis.
Requirement Guidelines Methods Type Target
@@
T9.1: System Definition ISO/IEC 42001 (§§ 4.3, 6.1.2); ISO/IEC 23894 § 5.3 M9.1 Description of AI system – define intended purpose for Art 6 mapping and record system details C P
T9.2: Risk identification ISO/IEC 42001 (§§ 6.1.2, 7.5); ISO/IEC 23894 § 5.4; ISO 31000; ISO 31010; ISO/IEC 5259; ISO 8000; ISO/IEC TR 24027 M9.2 Structured risk identification feeding both Art. 6 classification and Art. 9 risk management C P
@@
T9.3: Risk analysis ISO/IEC 23894; ISO/IEC 42001; ISO 31000; ISO 31010; M9.3 Core classification step – apply AI Act Art. 6: Annex I & Annex III gates based on structured analysis C P
@@
T9.4: Risk evaluation & prioritisation ISO/IEC 23894; ISO/IEC 42001; ISO 31000; ISO 31010 M9.4 Finalise classification decision and justification recording C P
@@
T9.5: Risk control measures ISO/IEC 23894 §§ 5.5–5.7; ISO/IEC 42001; ISO 31000; ISO/IEC 24029-1/-2; ISO/IEC TR 24027; ISO/IEC 5259; ISO 8000; ISO/IEC 27001; 27005 M9.5 Risk control strategy documented and aligned with AI Act essential requirements C FP
T9.6: PMM: Planning ISO/IEC 42001 (§§ 8.4, 9.1, 7.5); ISO/IEC 23894 (§§ 5.3, 5.7); ISO 31000; ISO/IEC 24970; ISO/IEC 27001; 27005 M9.6 PMM system established with structured monitoring requirements C P
@@
T9.7: PMM: Data collection & analysis ISO/IEC 42001 (§§ 9.1, 9.3, 7.5); ISO/IEC 23894 (§§ 5.5–5.7); ISO 31000; ISO/IEC 24970; ISO/IEC 27001; 27005 M9.7 Continuous assessment and early detection of anomalies or new hazards T FP
@@
T9.8: PMM: Review & CAPA ISO/IEC 42001 (§§ 8.3, 10, 7.5); ISO/IEC 23894 (§§ 5.6, 5.8); ISO 31000 M9.8 CAPA implemented, system updated, continuous compliance maintained C P
T9.9: Incident reporting ISO/IEC 42001 (§§ 8.4, 9.1, 7.5); ISO/IEC 23894 (§§ 5.7, 5.8); ISO/IEC 24970; ISO/IEC 27035 M9.9 Compliant incident reporting + full documentation + integration of learnings C P

### v2

@@ 3.1. R1: Human Agency and Oversight
Article 14 of the AI Act establishes the principle of human agency and oversight as a cornerstone of trustworthy AI. It requires that AI systems be designed and implemented so that they operate under meaningful human control and that ultimate responsibility remains with human operators. In practice, this means that individuals supervising the system must understand its capabilities and limitations, be able to intervene when necessary, and override its decisions to prevent or mitigate risks.
Requirement References Methods Type Target
R1.1: Human in the loop Sheridan et al. (sheridan1992telerobotics) M1.1: Manual override — Confirm that a human operator can take control of the AI system at any time during operation. C P, FP
Sheridan et al. (sheridan1992telerobotics) M1.2: Supervisory control — Assess whether human operators are able to supervise the AI system and intervene when necessary. C P, FP
@@
R1.2: User Autonomy GDPR (Art.7); Andreotta et al. (andreotta2022ai) M1.4: Informed consent — Check that users are provided with clear information about AI usage and are able to give or withhold consent. C FP
ISACA manual (§ 3.2) M1.5: User preferences — Review whether the system allows adaptation of its behaviour based on user-set preferences. C FP
ISACA manual (§ 3.2) M1.6: Opt-in/out — Confirm that users can enable or disable AI functionalities according to their choices. C FP
R1.3: Oversight Mechanisms ISACA manual (§ 2.11.1) M1.7: Audit trails — Ensure that system actions and decisions are recorded in tamper-evident logs for traceability. C P
@@ 3.2. R2: Technical Robustness and Safety
Requirement References Methods Type Target
R2.1: Resilience & Reliability ISACA manual (§ 1.4.4) M2.1: Stress testing — Assess whether the system has been exposed to extreme loads or edge cases to identify potential failure points. T P
ISACA manual (§ 2.8.6) M2.2: Drift detection — Confirm that mechanisms are in place to monitor data distribution changes and evaluate behaviour under such shifts. T D, P
@@
ISACA manual (§ 2.9.4) M2.4: Fail-over architecture — Ensure redundant components exist to maintain functionality in case of failure. C P, FP
R2.2: Robustness to Attacks Brundage et al. (Brundage2020); ISACA manual (§ 2.13.3) M2.5: Red teaming — Examine whether adversarial testing is conducted to uncover vulnerabilities. T P, FP
R2.3: Fallback & Fail-safe HLEG Guidelines M2.6: Human fallback — Check that the system includes a mechanism to transfer control to a human in the event of failure. C P, FP
R2.4: Accuracy & Uncertainty ISACA manual (§ 1.4.3) M2.7: Model calibration — Review whether confidence scores are appropriately aligned with model performance. T M
@@ 3.3. R3: Privacy and Data Governance
Article 10 of the AI Act establishes specific obligations concerning data governance and data quality for high-risk AI systems, requiring that datasets be relevant, representative, free of errors, and complete to the extent possible. These obligations are AI-specific and apply irrespective of whether personal data is involved. Where high-risk AI systems rely on personal data, the operational interpretation and verification of Article 10 requirements necessarily draw on established data protection mechanisms defined under the GDPR, in particular Articles 5, 25, and 32, which address purpose limitation, data protection by design, and security of processing. In such cases, GDPR concepts and safeguards provide concrete, widely operationalised mechanisms through which certain data governance obligations of the AI Act can be implemented and assessed in practice. Accordingly, compliance with Article 10 typically involves a combination of organisational and technical measures, such as encryption, pseudonymisation, access controls, and accountability procedures, which are already well established in data protection practice. Effective data governance thus supports compliance with AI Act requirements while also contributing to the technical robustness and trustworthiness of AI systems across their lifecycle.
Requirement References Methods Type Target
@@
GDPR (Art. 32.1.a); ISACA manual (§ 2.7.3) M3.5: Encryption — Check that data is encrypted in storage and transmission. T D, P
R3.4: Data Provenance ISACA manual (§ 2.7.5) M3.6: Lineage tracking — Confirm that data transformations are logged to ensure traceability. C D, P
ISACA manual (§ 3.1.2) M3.7: Metadata versioning — Verify dataset versions and schema changes are tracked. C D, P
@@ 3.4. R4: Transparency
Transparency is one of the most fundamental requirements for trustworthy and accountable AI. Under the AI Act, specifically in Art. 13, AI systems must be designed and documented in a way that allows users, regulators, and auditors to understand their capabilities, limitations, and decision-making logic. Transparency encompasses both technical and communicative aspects: it includes model explainability, data and process traceability, disclosure of AI usage to end users, and interpretability of outputs in accessible language. Together, these mechanisms ensure that AI operations remain observable and interpretable across all stages of their lifecycle, thereby fostering accountability and enabling meaningful human oversight.
Requirement References Methods Type Target
R4.1: Explainability ISACA manual (§ 3.2) M4.1: Local interpretability — Verify that the system provides human-understandable explanations of model outputs using recognised local explanation techniques (e.g., feature attribution or surrogate models). T D, M, FP
Mundhenk et al. (mundhenk2020) M4.2: Saliency maps — Check if visualisation tools highlight relevant input regions. T D, M, FP
R4.2: Model & Data Traceability GPAI CoP (Sec. Transparency) M4.3: Model cards — Ensure that standardised documentation is available and maintained. C M, P
@@
R4.3: Disclosure of AI Use ISO/IEC 22989 (§5.15) M4.5: AI identity notices — Verify that users are clearly informed when interacting with an AI system. C FP
GPAI CoP; ISACA manual (§ 3.2) M4.6: Disclaimers — Check whether limitations and assumptions are transparently disclosed. C FP
ISO/IEC 23894 (§6.1) M4.7: Risk communication — Assess whether the system communicates known risks and uncertainties. C FP
R4.4: Interpretability for End Users ISO/IEC 22989 (§5.15) M4.8: Understandable language — Confirm that technical terms are explained using plain language. C FP
@@ 3.5. R5: Diversity, Non-discrimination and Fairness
Diversity, non-discrimination, and fairness are central to ensuring that AI systems respect fundamental rights and avoid reinforcing structural biases. Articles 10, and 27 of the AI Act require that systems be designed, trained, and deployed in a way that prevents discriminatory outcomes and ensures inclusive access. This includes assessing data representativeness, applying fairness metrics across protected groups, and incorporating participatory design practices that reflect societal diversity. Fairness thus spans both technical and organisational domains, linking data quality, model evaluation, and stakeholder engagement to the ethical and legal accountability of AI systems.
Requirement References Methods Type Target
R5.1: Fairness in Data ISACA manual (§ 3.2) M5.1: Bias detection — Evaluate whether datasets are examined for representation gaps. T D, P
ISACA manual (§ 3.2) M5.2: Sampling strategies — Confirm that sampling ensures balanced representation. C D, P
R5.2: Fairness in Algorithms ISACA manual (§ 3.2) M5.3: Protected-class fairness metrics — Verify that fairness metrics are applied across protected groups. T D, M, P
R5.3: Inclusive Design WCAG 2.2 M5.4: Accessibility — Check that the system supports assistive technologies. C FP
R5.4: Stakeholder Engagement Schuler et al. (schuler1993participatory) M5.5: Participatory design — Assess whether affected users and communities are involved in design and testing. C FP, P
@@ 3.6. R6: Societal and Environmental Well-being
Requirement References Methods Type Target
R6.1: Environmental Impact UNESCO (Rec. 84); ISACA manual (§ 1.17.7) M6.1: Energy-use tracking — Verify that the system monitors computational energy consumption and carbon footprint. T P, FP
R6.2: Ethical Alignment Leslie et al. (CoE_AI_2020) M6.2: Rights-based system — Assess whether system design aligns with fundamental rights. C P
@@ 3.7. R7: Accountability
Requirement References Methods Type Target
R7.1: Responsibility Assignment ISACA manual (§ 3.2) M7.1: RASCI charts — Confirm that responsibility matrices define roles for each lifecycle activity. C P
R7.2: Auditability ISO/IEC 42001:2023 (Ann.B.6.2.8); ISACA manual (§ 2.11.1) M7.2: Logging — Verify that detailed, tamper-evident logs of system operations are maintained and reviewed. C P
@@
R7.4: Documentation Integrity ISO/IEC 42001:2023 (Ann.B.6.2.5); ISACA manual (§ 2.11.1) M7.5: Version-controlled records — Assess whether documentation is versioned and traceable over time. C P
@@ 3.8. R8: Quality Management
Article 17 of the AI Act requires providers of high-risk AI systems to establish, implement, document, and maintain a quality management system (QMS). The QMS ensures that the organisation applies consistent procedures for design, testing, data management, and post-market monitoring. It provides a structured approach for achieving and maintaining conformity with the requirements of the AI Act. Quality management thereby supports traceability, continuous improvement, and long-term compliance across the AI system lifecycle.
Requirement References Methods Type Target
@@ 3.9. R9: Risk Management
Article 9 of the AI Act establishes the obligation to implement a continuous and systematic risk management process throughout the lifecycle of high-risk AI systems. This process covers the identification, analysis, evaluation, and control of risks, as well as post-market monitoring and review. It must ensure that residual risks are acceptable relative to the intended use and that appropriate mitigation measures are applied. The methodology used here aligns with a set of horizontal, cross-sectoral standards, including ISO/IEC 23894 (AI risk management), ISO/IEC 42001 (AI management system), ISO 31000 and ISO 31010 (General risk management principles and techniques), ISO/IEC 22989 (AI concepts and terminology) as well as data- and robustness-related standards such as ISO/IEC 5259 series, ISO 8000 series, ISO/IEC TR 24027, ISO/IEC 24029-1/-2, and ISO/IEC 24970. In the future, when harmonised standards under the AI Act, particularly prEN 18228 (AI Risk Management) and prEN 18286 (AI Quality Management System) will be finalised, they will provide the solid structure for conformity verification and presumption of conformity once a system is categorised as high-risk.
@@
However, it is essential to underline that none of these standards, both international ISO/IEC standards and future harmonised European standards, determine whether an AI system is classified as high-risk under Article 6. Classification is governed exclusively by the legal criteria of the AI Act, particularly in Annex III. Standards are used in this report solely to support structured analysis, documentation and evidence generation, and they become fully applicable only after an AI system has been classified as high-risk. In addition, the methodology deliberately excludes sector-specific or application-specific standards, such as functional-safety standards (e.g., IEC 61508 or ISO 13849), medical-device standards (e.g., ISO 14971) or automotive safety standards (e.g., ISO 26262). These frameworks are highly relevant within their respective regulatory regimes but are not part of the horizontal, cross-sector risk-classification process defined by the AI Act and therefore are not included in the present analysis.
Requirement References Methods Type Target
@@
T9.1: System Definition ISO/IEC 42001 (§§ 4.3, 6.1.2); ISO/IEC 23894 (§ 6.3); ISO/IEC 22989 M9.1 Description of AI system – define intended purpose for Art 6 mapping and record system details C P
T9.2: Risk identification ISO/IEC 42001 (§§ 6.1.2, 7.5); ISO/IEC 23894 (§§ 5.4, 6.4); ISO 31000; ISO 31010; ISO/IEC 5259-1; ISO 8000-8; ISO/IEC TR 24027 M9.2 Structured risk identification relevant to classification (Art. 6) and risk management (Art. 9) C P
@@
T9.3: Risk analysis ISO/IEC 23894 (§§ 6.4, 6.5); ISO/IEC 42001; ISO 31000; ISO 31010; ISO/IEC 24029-2 M9.3 Core classification step – apply AI Act Art. 6: Annex I & Annex III gates using structured analysis C P
@@
T9.4: Risk evaluation & prioritisation ISO/IEC 23894 (§§ 6.4, 6.5); ISO/IEC 42001; ISO 31000; ISO 31010 M9.4 Finalize classification decision;record justification and prioritise risks C P
@@
T9.5: Risk control considerations ISO/IEC 23894 (§§ 5.5–5.7); ISO/IEC 42001; ISO 31000; ISO/IEC 24029-1/-2; ISO/IEC TR 24027; ISO/IEC 5259; ISO 8000; ISO/IEC 27001; ISO/IEC 27005 M9.5 Outline potential risk areas aligned with AI Act essential requirements C FP
T9.6: PMM: Planning ISO/IEC 42001 (§§ 8.4, 9.1, 7.5); ISO/IEC 23894 (§§ 5.3, 5.7); ISO 31000; ISO/IEC 24970; ISO/IEC 27001; ISO/IEC 27005 M9.6 Establishing PMM plan with monitoring, logging and reporting C P
@@
T9.7: PMM: Data collection & analysis ISO/IEC 42001 (§§ 9.1, 9.3, 7.5); ISO/IEC 23894 (§§ 5.5–5.7); ISO 31000; ISO/IEC 24970; ISO/IEC 27001; ISO/IEC 27005 M9.7 Continuous verification and early detection of anomalies or new hazards T FP
@@
T9.8: PMM: Review & CAPA ISO/IEC 42001 (§§ 8.3, 10, 7.5); ISO/IEC 23894 (§§ 5.6, 5.8); ISO 31000 M9.8 Perform CAPA, update system, maintain continuous compliance C P
T9.9: Incident reporting ISO/IEC 42001 (§§ 8.4, 9.1, 7.5); ISO/IEC 23894 (§§ 5.7, 5.8); ISO/IEC 24970; ISO/IEC 27035 M9.9 Compliant incident reporting; documentation; integrate lessons learned C P